Evaluate the following limits.

1. limX→+∞ 5X+6

2. limX→−∞ e4X+7

3.limX→−∞(0.2)X


1
Expert's answer
2022-03-20T06:40:18-0400

"1) \lim\limits _{x\to+\infty}(5x+6)=+\infty\\
2) \lim\limits _{x\to-\infty}(e^{4x}+7)=0+7=7\\
3) \lim\limits _{x\to-\infty}(0.2)^x=+\infty"

"y=e^{4x}"


"y=0.2^x"